OpenStudy (anonymous):

Compute the modulus and argument of each complex number. a. 1 + i b. 4 – 4i c. 2 + 5i d. 2[cos (2pi/3) + i sin (2pi/3)]

OpenStudy (anonymous):

The modulus is sqrt(a^2 + b^2) I will do one of them, you can finish the rest Modulus of 4-4i....a = 4, b = -4 Modulus = sqrt (16 + 16) = sqrt (32)

OpenStudy (anonymous):

thanks! how would you figure out the argument though?

OpenStudy (anonymous):

@Easyaspi314

OpenStudy (anonymous):

 and then just use tan